# A PBN-RL-XAI Framework for Discovering a “Hit-and-Run” Therapeutic Strategy in Melanoma

### v4

@@ Abstract
Index Terms:
Probabilistic Boolean Networks, Melanoma, Immunotherapy Resistance, Systems Biology, Explainable AI.
@@ I Introduction
The challenge of understanding resistance networks is compounded by their dynamic and stochastic nature. Gene regulatory networks exhibit complex emergent behaviors, including multistability, where identical genetic backgrounds can give rise to distinct cellular phenotypes depending on the network’s dynamical state. These systems can become trapped in stable attractor states corresponding to different cellular phenotypes, such as therapy resistance [3]. Identifying effective therapeutic interventions requires understanding not only the static architecture of these networks but also their dynamic properties and control mechanisms.
@@
Probabilistic Boolean Networks (PBNs) offer a powerful framework for modeling gene regulatory systems, combining computational tractability with biological realism through stochastic dynamics [5]. PBNs have successfully modeled various biological phenomena, including cell cycle regulation and cancer progression [6]. Concurrently, advances in reinforcement learning (RL) have opened new avenues for discovering optimal control strategies in complex dynamic systems [7]. Unlike traditional optimization approaches requiring explicit system knowledge, RL agents learn optimal policies through environment interaction, showing promise in biological control applications [8].
Here, we present a novel computational framework integrating PBN modeling with reinforcement learning to systematically discover optimal therapeutic interventions for overcoming innate anti-PD-1 resistance in melanoma. We constructed dynamic PBN models from patient transcriptomic data to capture the regulatory logic governing therapy response, then employed RL agents to identify multi-step intervention strategies. Using explainable artificial intelligence, we mechanistically interpreted the learned control policies to understand their biological basis.
Our analysis reveals that innate resistance corresponds to a fundamental system-level transition from a plastic, multi-stable network state to a rigid, canalized landscape dominated by a JUN/LOXL2 regulatory axis. Most significantly, our RL approach discovered a non-obvious ”hit-and-run” therapeutic strategy: a precisely timed, 4-step temporary inhibition of LOXL2 achieving 93.45% success rate in silico. Explainable AI analysis demonstrated that this transient intervention effectively erases the molecular vulnerability signature driving resistance, allowing the network’s intrinsic dynamics to complete the therapeutic transition without sustained intervention.
This work provides the first application of integrated PBN-RL-explainable AI(XAI) methodology to cancer immunotherapy resistance, offering a generalizable framework for discovering time-dependent therapeutic strategies in complex biological systems. Our work is organized as follows. Section II details our methodology, from data acquisition and network inference to the reinforcement learning and XAI frameworks. Section III presents the results, including the characterization of the resistant network and the discovery and validation of optimal control strategies. Finally, Section IV discusses the implications of our findings and concludes the paper.
@@ II-A Data Acquisition and Binarization
We utilized the landmark RNA-sequencing dataset from Hugo et al. [1], publicly available from the Gene Expression Omnibus (accession: GSE78220). Pre-treatment tumor biopsies were stratified into “Responders” (n=15), defined as patients with a Complete or Partial Response, and “Non-Responders” (n=12), defined as patients with Progressive Disease. One on-treatment sample was excluded to ensure the analysis focused exclusively on the innate resistance state. For biological interpretation, gene identifiers in the raw data were mapped to official HUGO Gene Symbols.
To prepare the data for Boolean modeling, the continuous gene expression values were converted into discrete binary states of 0 (OFF/low activity) or 1 (ON’/high activity). The analysis began with the raw count matrix, which was first filtered to include only the core genes. To stabilize the variance across the range of mean expression values, a Variance Stabilizing Transformation (VST)[9] was applied to the filtered count data using the DESeq2 R package[10]. The resulting normalized matrix was then separated into two sub-matrices, one for each cohort.
For each gene within each cohort’s VST-normalized matrix, we employed a Gaussian Mixture Model (GMM)[11] to determine a data-driven binarization threshold. This model-based approach was chosen over simpler methods like a median split because it provides a more principled cutoff. GMM assumes that the observed expression values are generated from two underlying distributions representing ’low’ and ’high’ activity states, and it mathematically identifies the threshold that best separates these putative states rather than imposing an arbitrary split.
Mathematically, the probability density of an expression value $x$ for a given gene is modeled as a weighted sum of two Gaussian components:
@@
where $\pi_{1}$ and $\pi_{2}$ are the mixing coefficients that sum to 1, $\mathcal{N}(x|\mu,\sigma^{2})$ is the Gaussian probability density function, and $(\mu_{\text{low}},\mu_{\text{high}},\sigma^{2})$ are the means and common variance of the two components, respectively. These parameters were fitted to the expression data for each gene using the mclust R package. The binarization threshold $T$ was then defined as the midpoint between the means of the two fitted components, $T=(\mu_{\text{low}}+\mu_{\text{high}})/2$. Samples with an expression value greater than this threshold were assigned a state of 1, while those with values less than or equal to the threshold were assigned a state of 0. This process yielded two final binarized matrices that served as the direct inputs for the subsequent network inference.
@@ II-B Introduction of PBN
To capture the complex and stochastic nature of gene regulatory networks (GRNs), we employ the Probabilistic Boolean Network (PBN) framework [5]. A PBN is a discrete-time dynamical system that models a set of genes as binary variables, where the state of each gene (ON=1, OFF=0) is updated at each time step according to a set of probabilistic rules.
A PBN is formally defined by a set of $n$ genes, $G=\{g_{1},g_{2},\ldots,g_{n}\}$. The state of the network at time $t$ is given by a binary vector $\mathbf{x}(t)=(x_{1}(t),x_{2}(t),\ldots,x_{n}(t))$, where $x_{i}(t)\in\{0,1\}$ is the state of gene $g_{i}$.
For each gene $g_{i}$, there exists a set of $k_{i}$ possible Boolean predictor functions, $F_{i}=\{f_{1}^{(i)},f_{2}^{(i)},\ldots,f_{k_{i}}^{(i)}\}$. Each function $f_{j}^{(i)}:\{0,1\}^{n}\to\{0,1\}$ determines the next state of gene $g_{i}$ based on the current state of all genes in the network. At each time step $t$, a single function $f_{j}^{(i)}$ is selected from the set $F_{i}$ with a corresponding probability $c_{j}^{(i)}$, where:
$P(f=f_{j}^{(i)})=c_{j}^{(i)}\quad\text{and}\quad\sum_{j=1}^{k_{i}}c_{j}^{(i)}=1$ (2)
The state of gene $g_{i}$ at the next time step, $t+1$, is then determined by the selected function: $x_{i}(t+1)=f_{j}^{(i)}(\mathbf{x}(t))$. The selection of a predictor function for each gene is an independent event, defining the transition probabilities between the $2^{n}$ possible states of the network.
The long-term behavior of a PBN is characterized by its attractors. An attractor is a set of states from which the network cannot escape once entered. Mathematically, a non-empty set of states $A$ is an attractor if it is a minimal set for which the probability of transitioning from any state $\mathbf{x}\in A$ to any state $\mathbf{y}\notin A$ is zero. These attractors can be a single steady state (a point attractor) or a set of states visited cyclically (a cycle attractor). In the context of systems biology, attractors represent the stable, long-term phenotypes of the cell, such as proliferation, apoptosis, or in our case, a therapy-resistant state[12].
To quantify the regulatory impact of one gene on another within the network, we utilize the influence of gene $g_{j}$ on gene $g_{i}$. This influence is based on the concept of the sensitivity of the Boolean predictor functions. The influence of $g_{j}$ on $g_{i}$ with respect to a specific function $f_{k}^{(i)}$ is the probability that flipping the state of $g_{j}$ will flip the output of $f_{k}^{(i)}$. The total influence, $I(g_{j}\to g_{i})$, is the expected sensitivity, averaged over all possible predictor functions for gene $g_{i}$:
@@
where $S(f_{k}^{(i)},g_{j})$ is the sensitivity of function $f_{k}^{(i)}$ to gene $g_{j}$. The influence matrix, composed of these values, provides a quantitative map of the network’s regulatory structure.
@@ II-C PKN-constrained PBN construction
The inference of Probabilistic Boolean Networks (PBNs) from data is an NP-hard problem, where the search space for regulatory logic grows double-exponentially with the number of inputs to a given gene, making the analysis of large-scale networks computationally infeasible. Therefore, a critical preliminary step is to constrain the problem space by identifying a core set of the most biologically relevant and dynamically active genes. To this end, we implemented a multi-stage procedure to construct a high-confidence, context-specific Prior Knowledge Network (PKN) that serves as the basis for our tractable PBN model.
II-C1 Initial network foundation
The process began with a broad, literature-based network that was systematically refined into a compact, data-driven core network. The initial network foundation was based on the Kyoto Encyclopedia of Genes and Genomes (KEGG) pathway hsa05235, ”PD-L1 expression and PD-1 checkpoint pathway in cancer,” chosen for its direct relevance to the therapeutic mechanism under investigation. The pathway data was obtained in KGML format, and all constituent gene-gene interactions were parsed and mapped to their corresponding HUGO Gene Symbols.
II-C2 Hybrid gene selection strategy
To distill this foundational network into a focused set of key regulators, we devised a hybrid selection strategy combining mandatory inclusion of expert-defined markers with a data-driven ranking of all other genes. This approach guaranteed the inclusion of five core genes from the Innate anti-PD-1 Resistance (IPRES) signature: AXL, ROR2, WNT5A, LOXL2, and TAGLN. The remaining positions in the core gene list were filled by ranking genes based on a composite score derived from two metrics.
The first metric, a Differential Expression Score ($S_{DE}$), was calculated as the negative base-10 logarithm of the Benjamini-Hochberg adjusted p-value ($-\log_{10}(p_{adj})$)[13] from a differential expression analysis[14] between the responder and non-responder cohorts. The second metric, a Network Influence Score ($S_{Net}$), was calculated from a data-driven gene regulatory network inferred by dynGENIE3[15], capturing a gene’s overall importance as both a regulator and a target.
For the final PKN assembly, both the $S_{DE}$ and $S_{Net}$ scores were min-max normalized to a [0, 1] scale. The final data-driven score for each gene was calculated using the equally weighted formula:
$S_{final}=(0.5\cdot S_{DE,norm})+(0.5\cdot S_{Net,norm})$ (4)
The final core gene list was created by taking the union of the five mandatory IPRES genes and the top 12 genes from this data-driven ranking. The final PKN was then constructed by extracting the subnetwork from the dynGENIE3-inferred weight matrix corresponding only to the interactions among these selected core genes.
II-C3 PBN inference methodology
With the binarized data for the Responder and Non-Responder cohorts, we inferred two distinct PBN models. The inference strategy was designed to identify the most informative regulatory logic for each gene while mitigating the risk of overfitting the model to noise in the data.
The process began by constraining the search space for potential regulators for each target gene. This is a critical step for computational tractability, as testing all possible gene combinations is infeasible. We used the global gene-gene interaction rankings previously generated by dynGENIE3 as a data-driven prior. For each target gene, we selected its top-ranked regulators from this list to form a candidate pool, limiting the pool size to a maximum of 10 genes to balance biological comprehensiveness with computational efficiency.
A key challenge in network inference is selecting the correct number of input regulators, $k$, for each gene’s predictive function. A model that is too simple (small $k$) may underfit and miss critical interactions, while one that is too complex (large $k$) may overfit by modeling noise. To navigate this trade-off, we implemented a data-driven selection criterion based on maximizing the Mutual Information (MI). For each target gene, our algorithm iteratively tested regulator sets of increasing size (up to a maximum of $k_{max}=4$). At each value of $k$, the algorithm identified the single best-predicting Boolean function from all $2^{2^{k}}$ possibilities and calculated its MI score with the target gene’s observed data. The optimal number of regulators, $k^{*}$, was chosen as the value of $k$ that yielded the function with the maximum MI score.
With an optimal regulator set of size $k^{*}$ identified for a given target gene, its probabilistic Boolean update rule was defined. To account for biological stochasticity and measurement noise, we constructed a probabilistic ensemble of the most likely functions rather than selecting only the single best predictor. All $2^{2^{k^{*}}}$ possible Boolean functions were scored using their normalized Mutual Information (MI)[16] value, and the top $N=4$ functions were selected to form the function set $\{f_{1},f_{2},\ldots,f_{N}\}$ for that gene. The probability $P(f_{i})$ of each function was then assigned proportionally to its MI score, as follows:
@@ II-C3 PBN inference methodology
where $MI(f_{i})$ is the Mutual Information score of function $f_{i}$. An alternative ranking approach using the Coefficient of Determination (CoD)[17] was considered; however, it was found to be impractical as it produced negative values for some functions in our dataset, making MI a more robust metric for this application.
This entire procedure was performed independently for every gene in both the responder and non-responder datasets, yielding the two final, context-specific PBN models used for all subsequent analysis.
@@ II-D Reinforcement Learning for Optimal Dynamic Control
To discover an optimal intervention strategy in a complex biological network, we framed the control of the resistant-state PBN as a reinforcement learning (RL) problem. The state space of the PBN is vast ($2^{n}$ states), rendering methods based on exhaustive search computationally intractable. Furthermore, the inherent stochasticity of the network means a fixed sequence of interventions does not guarantee a specific outcome. RL is well-suited for finding optimal policies in such large, stochastic environments. To formulate the control problem, we utilized the gym-PBN software package, a toolkit designed for the analysis and RL control of Probabilistic Boolean Networks within the Gymnasium framework [18].
We formally model the control problem as a Markov Decision Process (MDP) [19]. The state space, $\mathcal{S}$, comprises the $2^{n}$ possible binary states, $\mathbf{x}$, of the PBN. The action space, $\mathcal{A}$, is a set of discrete actions including “Do Nothing” and flipping the state of any single gene. The transition probabilities, $P(\mathbf{x}_{t+1}|\mathbf{x}_{t},a_{t})$, are implicitly defined by the PBN’s update rules under a given action. To model the clinical time constraint, we set the maximum steps to be 15. Finally, the reward function, $R$, was designed to guide the agent: a large positive reward ($+100$) is given for reaching the target sensitive state, and a penalty ($-5$) for ending an episode in a known resistant attractor. The action cost is set to 00 so that the agent is not prevented from imposing actions during the 15 steps horizon. The agent’s goal is to learn a policy, $\pi(a_{t}|\mathbf{x}_{t})$, that maximizes the expected cumulative reward.
We chose Proximal Policy Optimization (PPO) as our learning algorithm due to its state-of-the-art performance and stability [20]. PPO optimizes the clipped surrogate objective:
@@
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|\mathbf{x}_{t})}{\pi_{\theta_{old}}(a_{t}|\mathbf{x}_{t})}$ is the probability ratio between new and old policies, $A_{t}$ is the advantage estimate, and $\epsilon$ is the clipping parameter. This clipping mechanism prevents large, destabilizing policy updates by constraining $r_{t}(\theta)\in[1-\epsilon,1+\epsilon]$. We utilized the PPO agent from Stable-baselines3 [21] with default hyperparameters ($\gamma=0.99$, $\lambda=0.95$, $\epsilon=0.2$) and trained for 500,000 timesteps.
@@ II-E1 Optimal “Hit-and-Run” Strategy Identification
To identify a more clinically translatable strategy, we also sought to find an optimal transient intervention. We designed a sequential experiment to test a “hit-and-run” therapeutic model, where the rationale is that a brief, initial perturbation may be sufficient to destabilize the resistant state, allowing the network’s own dynamics to resolve towards a sensitive phenotype without sustained intervention. This experiment was conducted in two phases. First, in the priming phase, we simulated a temporary inhibition of a key target gene (LOXL2, MAPK3, or JUN) by clamping its state to 0 for a specific duration, ranging from 1 to 5 simulation steps. Second, in the control phase, the pre-trained PPO agent was given control of the network from its now-primed state. The agent’s success in steering the network to the target sensitive state within the remainder of the 15-step episode was recorded to identify the most effective transient intervention.
@@ II-E2 Mechanistic Policy Explanation via SHAP
To gain biological insight into our agent’s decision-making strategy, we employed SHAP (SHapley Additive exPlanations) [22], a game-theoretic explainable AI method that quantifies each feature’s contribution to model predictions by computing average marginal contributions across all possible feature combinations. We applied SHAP to analyze the agent’s policy in two key scenarios: (1) identifying the context-dependent ”vulnerability signature” of genes that triggered specific interventions, and (2) assessing how transient priming interventions altered this signature. Additionally, we generated SHAP trajectory visualizations to track the evolution of feature importance over entire control episodes, enabling dynamic interpretation of the agent’s reasoning process.

### v5

@@ I Introduction
While existing computational approaches for therapeutic discovery have explored various methodologies—including Boolean network control [4], dynamical systems optimization [5], and machine learning-based drug repurposing [6]—these methods typically focus on static network analysis or sustained interventions. Our approach uniquely integrates Probabilistic Boolean Networks (PBNs) with reinforcement learning (RL) to discover time-dependent, transient intervention strategies. PBNs combine computational tractability with biological realism through stochastic dynamics [7], while RL enables discovery of optimal control policies without requiring explicit system knowledge [8]. This integration allows systematic exploration of dynamic intervention strategies that existing static approaches cannot capture.
Here, we present this novel computational framework to systematically discover optimal therapeutic interventions for overcoming innate anti-PD-1 resistance in melanoma. We constructed dynamic PBN models from patient transcriptomic data to capture the regulatory logic governing therapy response, then employed RL agents to identify multi-step intervention strategies. Using explainable artificial intelligence (XAI), we mechanistically interpreted the learned control policies to understand their biological basis.
Our analysis reveals that innate resistance corresponds to a fundamental system-level transition from a plastic, multi-stable network state to a rigid, canalized landscape dominated by a JUN/LOXL2 regulatory axis. Most significantly, our RL approach discovered a non-obvious “hit-and-run” therapeutic strategy: a precisely timed, 4-step temporary inhibition of LOXL2 achieving 93.45% success rate in silico. This computational hypothesis requires rigorous experimental validation before clinical consideration, yet demonstrates the framework’s potential for discovering non-intuitive therapeutic strategies in complex biological systems.
@@ II-A Data Acquisition and Binarization
We utilized the landmark RNA-sequencing dataset from Hugo et al. [1], publicly available from the Gene Expression Omnibus[9] (accession: GSE78220). This dataset comprises transcriptomic profiles from 28 pre-treatment melanoma biopsies obtained from patients undergoing anti-PD-1 checkpoint inhibitor therapy (pembrolizumab or nivolumab) for metastatic melanoma. The cohort included patients aged 19-84 years (median 60 years) with 75% male representation and predominantly advanced disease (89% M1b/M1c stage). Approximately 50% had prior MAPK inhibitor treatment.
Pre-treatment tumor biopsies were stratified into “Responders” (n=15), defined as patients achieving Complete or Partial Response, and “Non-Responders” (n=13), defined as patients with Progressive Disease according to immune-related response criteria. One on-treatment sample was excluded to ensure analysis focused exclusively on the innate resistance state. Gene identifiers were mapped to official HUGO Gene Symbols for biological interpretation.
To prepare data for Boolean modeling, continuous gene expression values were converted into discrete binary states of 0 (OFF/low activity) or 1 (ON/high activity). Raw count data were filtered to include only core genes, then Variance Stabilizing Transformation (VST) [10] was applied using DESeq2 [11] to stabilize variance across expression ranges.
For each gene within each cohort’s VST-normalized matrix, we employed a Gaussian Mixture Model (GMM) [12] to determine data-driven binarization thresholds. The probability density of expression value $x$ was modeled as:
@@
where $\pi_{1}$ and $\pi_{2}$ are mixing coefficients summing to 1, and $(\mu_{\text{low}},\mu_{\text{high}},\sigma^{2})$ represent component means and variance. Parameters were fitted using mclust R package, with binarization threshold $T=(\mu_{\text{low}}+\mu_{\text{high}})/2$. This yielded two binarized matrices for subsequent network inference.
@@ II-B Introduction of PBN
To capture the complex and stochastic nature of gene regulatory networks (GRNs), we employ the Probabilistic Boolean Network (PBN) framework [7]. A PBN models $n$ genes as binary variables with state vector $\mathbf{x}(t)=(x_{1}(t),\ldots,x_{n}(t))$ where $x_{i}(t)\in\{0,1\}$ represents gene $g_{i}$’s state at time $t$.
For each gene $g_{i}$, there exists a set of $k_{i}$ Boolean predictor functions $F_{i}=\{f_{1}^{(i)},\ldots,f_{k_{i}}^{(i)}\}$. At each time step, function $f_{j}^{(i)}$ is selected with probability $c_{j}^{(i)}$ where $\sum_{j=1}^{k_{i}}c_{j}^{(i)}=1$, determining the next state: $x_{i}(t+1)=f_{j}^{(i)}(\mathbf{x}(t))$.
The long-term behavior is characterized by attractors—minimal sets of states $A$ where $P(\mathbf{x}(t+1)\in A|\mathbf{x}(t)\in A)=1$ and $P(\mathbf{x}(t+1)\notin A|\mathbf{x}(t)\in A)=0$. In systems biology, attractors represent stable cellular phenotypes such as therapy-resistant states [13].
To quantify regulatory relationships, we compute the influence of gene $g_{j}$ on $g_{i}$ as the expected sensitivity across all predictor functions:
@@
where $S(f_{k}^{(i)},g_{j})$ is the sensitivity of function $f_{k}^{(i)}$ to gene $g_{j}$. The resulting influence matrix provides a quantitative map of the network’s regulatory structure.
@@ II-C PKN-constrained PBN construction
PBN inference is computationally intractable for large networks due to double-exponential growth in search space. We developed a multi-stage procedure to construct a tractable, biologically-relevant 12-gene core network from melanoma anti-PD-1 resistance data.
We began with the KEGG pathway hsa05235 “PD-L1 expression and PD-1 checkpoint pathway” [14], directly relevant to anti-PD-1 therapy mechanisms. Our hybrid selection strategy mandatorily included five IPRES signature genes (AXL, ROR2, WNT5A, LOXL2, TAGLN) [1] known to drive therapy resistance through mesenchymal transition and angiogenesis pathways. The remaining seven genes were selected using a composite score combining differential expression significance ($S_{DE}$) calculated as $\log_{10}(p_{adj})$ from Benjamini-Hochberg adjusted p-values [15] from differential expression analysis [16], and network centrality ($S_{Net}$) from dynGENIE3 inference [17]:
$S_{final}=0.5\cdot S_{DE,norm}+0.5\cdot S_{Net,norm}$ (3)
The 12-gene limitation balances computational tractability ($2^{12}=4096$ states) with biological comprehensiveness. The selected genes span key resistance mechanisms: immune checkpoint regulation (PDCD1, CD274), mesenchymal transition (AXL, ROR2, WNT5A), extracellular matrix remodeling (LOXL2, TAGLN), and metabolic reprogramming. Prior studies demonstrate that core regulatory modules of 10-15 genes can capture essential cellular dynamics [7].
For PBN inference, we constrained potential regulators for each target gene using dynGENIE3 rankings, testing regulator sets of size $k=1$ to $k_{max}=4$. The optimal $k^{*}$ was selected by maximizing Mutual Information (MI) [18]. For each gene’s probabilistic update rule, we selected the top $N=4$ Boolean functions, with probabilities assigned proportionally:
@@
We also evaluated Coefficient of Determination (CoD) [19] as an alternative metric, but MI proved superior due to CoD’s susceptibility to negative values in our dataset.. This procedure generated two context-specific PBN models for Responder and Non-Responder cohorts, capturing distinct regulatory logic underlying therapy response patterns.
@@ II-D Reinforcement Learning for Optimal Dynamic Control
To discover optimal intervention strategies in the resistant-state PBN, we formulated the control problem as a Markov Decision Process (MDP) [20] using the gym-PBN framework [21]. The state space $\mathcal{S}$ comprises $2^{12}=4096$ binary network states, with action space $\mathcal{A}$ including “Do Nothing” and single-gene state flips. Transition probabilities $P(\mathbf{x}_{t+1}|\mathbf{x}_{t},a_{t})$ are defined by PBN update rules under interventions.
The reward function incentivizes therapeutic transitions: +100+100 for reaching sensitive states, 5-5 for resistant attractors, and 0 action cost to encourage active intervention within the 15-step clinical horizon. This design reflects the clinical priority of escaping resistance over minimizing intervention frequency.
We employed Proximal Policy Optimization (PPO) [22] for its proven stability in complex biological control tasks. PPO optimizes the clipped surrogate objective:
@@
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|\mathbf{x}_{t})}{\pi_{\theta_{old}}(a_{t}|\mathbf{x}_{t})}$ represents the policy ratio and $A_{t}$ the advantage estimate.
We employed standard PPO hyperparameters from the Stable-Baselines3  [23] implementation: discount factor $\gamma=0.99$ for long-term reward optimization, GAE parameter $\lambda=0.95$ to balance bias-variance in advantage estimation, and clipping parameter $\epsilon=0.2$ to prevent destabilizing policy updates while maintaining exploration. These hyperparameters were determined through preliminary experiments to ensure stable performance on our PBN model..
@@ II-E1 Optimal “Hit-and-Run” Strategy Identification
We designed a sequential experiment to test clinically translatable transient interventions, where brief perturbations destabilize resistant states without sustained therapy. The protocol consisted of two phases: a priming phase with temporary target gene inhibition (LOXL2, MAPK3, or JUN clamped to 0 for 1-5 steps), followed by a control phase where the pre-trained PPO agent guided the primed network toward sensitive states within the remaining 15-step episode. Success rates identified optimal transient intervention strategies.
@@ II-E2 Mechanistic Policy Explanation via SHAP
We employed SHAP (SHapley Additive exPlanations) [24] to analyze agent decision-making by quantifying each gene’s contribution to intervention choices through game-theoretic marginal contribution analysis. SHAP analysis identified context-dependent vulnerability signatures triggering specific interventions and assessed how transient priming altered these signatures. Trajectory visualizations tracked feature importance evolution across control episodes for dynamic interpretation of agent reasoning.